OpenStudy (anonymous):

what is the GCF of 20 and 30?

OpenStudy (alexandervonhumboldt2):

20=2*2*5 30=2*3*5

OpenStudy (alexandervonhumboldt2):

`The greatest common factor, or GCF, is the greatest factor that divides two numbers. To find the GCF of two numbers: List the prime factors of each number. Multiply those factors both numbers have in common. If there are no common prime factors, the GCF is 1.`

OpenStudy (anonymous):

its 10

OpenStudy (anonymous):

thank you i have another question what is the LCM 15 and 60

OpenStudy (anonymous):

umm 1 sec

OpenStudy (anonymous):

the answer is 60

OpenStudy (anonymous):

did you get it right?